# Investigating symptom duration using current status data: a case study of post-acute COVID-19 syndrome

### v2

@@ Results
After the removal of $284$ individuals due to missing covariate information, the analysis cohort consisted of $3434$ HCT study participants with positive SARS-CoV-2 tests prior to February $24$, $2023$. Of these, $1432$ ($41.7$%) responded to the survey. Characteristics of the analysis cohort are summarized in Table 3. The median age was $23$ years (IQR: 2020202036363636), and approximately $64$% of the participants reported female sex. Slightly over $33$% of participants reported at least one baseline comorbidity, and roughly $52$% were symptomatic at the time of testing. Most participants were not recently vaccinated. Survey respondents tended to be older than nonrespondents, with markedly higher prevalence of symptoms during the acute infection phase.
@@
        25 and under 597 (41.7%) 1385 (69.2%) 1982 (57.7%)
        26–40 361 (25.2%) 379 (18.9%) 740 (21.5%)
        41–60 375 (26.2%) 183 (9.1%) 558 (16.2%)
        61 and over 99 (6.9%) 55 (2.7%) 154 (4.5%)
Male 494 (34.5%) 746 (37.3%) 1240 (36.1%)
Symptoms, upper respiratory 1027 (71.7%) 685 (34.2%) 1712 (49.9%)
Symptoms, systemic 715 (49.9%) 473 (23.6%) 1188 (34.6%)
Symptoms, gastrointestinal 122 (8.5%) 107 (5.3%) 229 (6.7%)
Symptoms, fatigue 465 (32.5%) 324 (16.2%) 789 (23.0%)
Comorbidities, cardiometabolic 96 (6.7%) 49 (2.4%) 145 (4.2%)
Comorbidities, respiratory 147 (10.3%) 193 (9.6%) 340 (9.9%)
Comorbidities, immunological 29 (2.0%) 26 (1.3%) 55 (1.6%)
Comorbidities, seasonal allergy 433 (30.2%) 439 (21.9%) 872 (25.4%)
@@
         Under 3 months 307 (21.4%) 27 (1.3%) 334 (9.7%)
         3–6 months 204 (14.2%) 123 (6.1%) 327 (9.5%)
         More than 6 months 921 (64.3%) 1852 (92.5%) 2773 (80.8%)
@@
         PCR, over 30 Ct 273 (19.1%) 438 (21.9%) 711 (20.7%)
         PCR, under 30 Ct 545 (38.1%) 897 (44.8%) 1442 (42.0%)
         Rapid antigen test 614 (42.9%) 667 (33.3%) 1281 (37.3%)
@@
Figure 1 shows the estimated proportion of participants with persistent COVID-19 symptoms at times from 30 to 90 days following the positive test, along with the distribution of survey response times. From the extended CIR method, estimated symptom persistence proportions at $30$, $60$, and $90$ days after the positive test were $19.0$% (95% CI: 13.613.613.613.625.025.025.025.0), $13.0$% (9.49.49.49.416.716.716.716.7), and $6.9$% (2.92.92.92.910.910.910.910.9), respectively.
@@
Table 4 contains the results of the Cox proportional hazards regression, with hazard ratios larger than $1$ indicating faster symptom resolution. Participants identifying as male tended to experience faster symptom resolution, with an estimated hazard ratio for resolution of $1.21$ (1.041.041.041.041.401.401.401.40, $P=0.018$). Conversely, participants reporting fatigue at baseline and participants with history of seasonal allergies tended to have longer lasting symptoms, with estimated hazard ratios of $0.77$ (0.650.650.650.650.920.920.920.92, $P=0.004$) and 0.86 (0.740.740.740.741.001.001.001.00, $P=0.050$), respectively. In terms of viral load, participants with Ct values under $30$ and participants who tested positive via rapid antigen tests had more persistent symptoms compared to participants with Ct values over $30$, with estimated hazard ratios of $0.78$ (0.640.640.640.640.960.960.960.96) and $0.75$ (0.600.600.600.600.950.950.950.95), respectively ($P=0.033$).
@@
        26–40 0.922 (0.767, 1.109)
        41–60 0.921 (0.754, 1.124)
        61 and over 1.078 (0.785, 1.480)
Male 1.206 (1.039, 1.401) 0.014
Symptoms, upper respiratory 1.078 (0.885, 1.314) 0.453
Symptoms, systemic 0.967 (0.803, 1.163) 0.719
Symptoms, gastrointestinal 0.887 (0.679, 1.160) 0.382
Symptoms, fatigue 0.771 (0.645, 0.921) 0.004
Comorbidities, cardiometabolic 0.808 (0.599, 1.088) 0.160
Comorbidities, respiratory 0.868 (0.694, 1.085) 0.214
Comorbidities, immunological 0.839 (0.489, 1.440) 0.525
Comorbidities, seasonal allergy 0.859 (0.737, 1.000) 0.050
Time since last vaccination 0.813
@@
         3–6 months 0.953 (0.752, 1.209)
         More than 6 months 0.934 (0.759, 1.150)
Viral load 0.033
@@
         PCR, under 30 Ct 0.783 (0.637, 0.963)
         Rapid antigen test 0.754 (0.597, 0.953)
@@ Discussion
A retrospective study in a comparable university population found that $36$% of survey respondents reported symptoms lasting at least 28 days from the end of the recommended $10$-day isolation period [26]. Our results provide further evidence that, even in a predominantly young and vaccinated population, a substantial proportion of infected individuals are affected by persistent COVID-19 symptoms. Notably, because participants were asked only about persistent symptoms and not about new symptoms arising in the post-acute phase, our analysis may still represent an underestimate of the frequency of PASC, which may also be characterized by symptoms that differ from initial acute symptoms and arise after their resolution.
We also investigated potential risk factors for persistent symptoms using Cox proportional hazards regression, and found evidence that female sex, fatigue during the acute phase, history of seasonal allergies, and higher semi-quantitative viral load at time of testing are associated with slower resolution of symptoms. This is in line with previous studies that have identified female sex [37], severe fatigue during the acute infection phase [38], allergic rhinitis [39], and higher peak semi-quantitative viral load in the four weeks following diagnosis [40] as potential risk factors for PASC.
The validity of extended CIR hinges on the untestable assumption that survey response times and event indicators are conditionally independent given measured baseline covariates, which is more likely to be satisfied through the inclusion of potential risk factors as covariates in the analysis. If the survey response time depends on symptom duration even within strata defined by baseline covariates, the extended CIR procedure could be susceptible to bias. The development of sensitivity analysis methods for this setting is an important avenue of future research. Further work is also needed to better understand the practical implications of the poor boundary performance of CIR in order to provide a principled approach to choose appropriate values of $t_{0},t_{1}$ and $c_{0}$.
@@ Acknowledgements
We denote by $T$ the unobserved symptom duration after testing positive for SARS-CoV-2. For each study participant, we observe a vector $W\in\mathbb{R}^{p}$ of baseline covariates. In the context of current status sampling, the ideal data unit — that is, the data unit that would be observed given unlimited follow-up — consists of $W$, the survey response time $Y_{*}\in[b_{0},\infty)$, and the event indicator $\Delta_{*}\in\{0,1\}$. We denote by $c_{0}$ the maximum follow-up time of the study. In reality, we observe $Y:=\min(Y,c_{0})$ and $\Delta:=\mathbbm{1}(Y_{*}<c_{0})\Delta_{*}$. For individuals who have not responded to the survey by time $c_{0}$, we do not observe the event indicator. Thus, the observed data unit is $Z:=(W,Y,\Delta)$. In this setting, survey non-response corresponds to having failed to respond to the survey prior to time $c_{0}$. The available data consist of $n$ independent observations $Z_{1},Z_{2},\ldots,Z_{n}$ draw from distribution $P_{0}$, which is assumed to lie in a nonparametric model $\mathcal{M}$. Here and after, we use the subscript 0 to denote a functional of the true data-generating distribution $P_{0}$, e.g., $E_{0}\{f(W)\}:=E_{P_{0}}\{f(W)\}$. For simplicity of notation, in what follows we consider estimation of the cumulative distribution function $t\mapsto\psi_{0}(t):=P_{0}(T\leq t)$. (In the main text, $\psi$ was used to denote the survival function, which is simply equal to one minus the distribution function).
@@
Without making strong modeling assumptions, the target regression function $\theta_{0}(y_{0})$ is not pathwise differentiable and is therefore not estimable at $n^{\frac{1}{2}}$–rate. However, it is known to be monotone as a function of $y_{0}$ and can be viewed as the target of an isotonic regression procedure. Using isotonic regression to estimate $\theta_{0}(y_{0})$ allows for the application of established statistical theory [41] to characterize the large-sample properties of the resulting estimator. We outline this isotonic regression procedure here.
For a generic $P\in\mathcal{M}$, we first define the outcome regression function $\mu_{P}:(y,w)\mapsto E_{P}(\Delta\,|\,W=w,Y=y)$ and density ratio $g_{P}:(y,w)\mapsto\pi_{P}(y\,|\,w)/f_{P}(y)$ with $\pi_{P}$ the conditional density of $Y$ given $W$ and $f_{P}$ the marginal density of $Y$. These nuisance functions play a key role in accounting for the possible dependence between $\Delta_{*}$ and $Y_{*}$. We use $F_{P}$ to denote the marginal distribution function of $Y$. We use $F_{P}$ as a domain transformation function in the CIR procedure. While the CIR procedure could be performed without the use of this domain transformation, its use results in a procedure that is invariant to strictly increasing transformations of $Y$ [30]. In other words, changing the scale on which response times are measured will have no impact on the estimated survival function.
@@
with $Q_{P}$ the marginal distribution function of $W$ under sampling from $P$. Unlike $\theta_{0}(y_{0})$, the parameter $\Gamma_{0}(y_{0})$ is pathwise differentiable relative to the nonparametric model $\mathcal{M}$ and can therefore be estimated at $n^{\frac{1}{2}}$–rate [42]. Per the results of Westling et al. [30], the efficient influence function of $P\mapsto\Gamma_{P}(y_{0})$ relative to $\mathcal{M}$ evaluated at $P_{0}$ is given by
@@
Knowledge of this efficient influence functions enables standard debiasing approaches, such as the one-step correction [43], to be used in estimation of $\Gamma_{0}(y_{0})$. This yields a debiased, asymptotically linear estimator $\Gamma_{n}(y_{0})$ of $\Gamma_{0}(y_{0})$, which is used in the isotonic regression procedure to construct an estimator of $\theta_{0}(y_{0})$. The procedure is as follows.
@@
    Let $\mathcal{R}_{n}$ denote the set of unique values of $Y_{1},Y_{2},\ldots,Y_{n}$. For each $y_{0}\in\mathcal{R}_{n}$ satisfying $t_{0}\leq y\leq t_{1}$, compute the one-step debiased estimator $\Gamma_{n}(y_{0})$ of $\Gamma_{0}(y_{0})$, given by $\Gamma_{n}(y_{0}):=\Gamma_{P_{n}}(y_{0})+\frac{1}{n}\sum_{i=1}^{n}D^{*}_{P_{n},y_{0}}(Z_{i})$.
@@
    Compute $\theta_{n}(y_{0}):=\psi_{n}(\Phi_{n}(y_{0}))$ as an estimator of $\theta_{0}(y_{0})$.
If both $\mu_{n}$ and $g_{n}$ converge to $\mu_{0}$ and $g_{0}$, Theorem 2 of Westling et al. [30] establishes that, for any $y_{0}\in(b_{0},c_{0})$ with $F_{0}(y_{0})\in(0,1)$, $n^{\frac{1}{3}}\left\{\theta_{n}(y_{0})-\theta_{0}(y_{0})\right\}$ converges in distribution to $\tau_{0}(y_{0})^{\frac{1}{3}}\mathbb{W}$, where $\mathbb{W}$ follows a standard Chernoff distribution [31] and $\tau_{0}(y_{0}):=4\theta_{0}^{\prime}(y_{0})\kappa_{0}(y_{0})/f_{0}(y_{0})$ with

### v3

@@ Results
After the removal of $285$ individuals due to missing covariate information, the analysis cohort consisted of $3489$ HCT study participants with positive SARS-CoV-2 tests prior to February $24$, $2023$. Of these, $1434$ ($41.1$%) responded to the survey. Characteristics of the analysis cohort are summarized in Table 3. The median age was $23$ years (IQR: 2020202036363636), and $64$% of the participants reported female sex. Slightly over $33$% of participants reported at least one baseline comorbidity, and roughly $52$% were symptomatic at the time of testing. Most participants were not recently vaccinated. Survey respondents tended to be older than nonrespondents, with markedly higher prevalence of symptoms during the acute infection phase.
@@
        25 and under 597 (41.6%) 1436 (69.9%) 2033 (58.3%)
        26–40 362 (25.2%) 385 (18.7%) 747 (21.4%)
        41–60 376 (26.2%) 179 (8.7%) 555 (15.9%)
        61 and over 99 (6.9%) 55 (2.7%) 154 (4.4%)
Male 495 (34.5%) 762 (37.1%) 1257 (36.0%)
Symptoms, upper respiratory 1028 (71.7%) 709 (34.5%) 1737 (49.8%)
Symptoms, systemic 715 (49.9%) 480 (23.4%) 1195 (34.3%)
Symptoms, gastrointestinal 122 (8.5%) 113 (5.5%) 235 (6.7%)
Symptoms, fatigue 466 (32.5%) 331 (16.1%) 797 (22.8%)
Comorbidities, cardiometabolic 96 (6.7%) 50 (2.4%) 146 (4.2%)
Comorbidities, respiratory 147 (10.3%) 196 (9.5%) 343 (9.8%)
Comorbidities, immunological 29 (2.0%) 25 (1.2%) 54 (1.5%)
Comorbidities, seasonal allergy 434 (30.3%) 445 (21.7%) 879 (25.2%)
@@
         Under 3 months 309 (21.5%) 265 (12.9%) 574 (16.5%)
         3–6 months 202 (14.1%) 314 (15.3%) 516 (14.8%)
         More than 6 months 923 (64.4%) 1476 (71.8%) 2399 (68.8%)
@@
         PCR, over 30 Ct 273 (19.0%) 452 (22.0%) 725 (20.8%)
         PCR, under 30 Ct 545 (38.0%) 924 (45.0%) 1469 (42.1%)
         Rapid antigen test 616 (43.0%) 679 (33.0%) 1295 (37.1%)
@@
Figure 1 shows the estimated proportion of participants with persistent COVID-19 symptoms at times from 30 to 90 days following the positive test, along with the distribution of survey response times. From the extended CIR method, estimated symptom persistence proportions at $30$, $60$, and $90$ days after the positive test were $19.2$% (95% CI: 13.713.713.713.725.225.225.225.2), $12.9$% (9.49.49.49.416.516.516.516.5), and $6.8$% (2.82.82.82.810.910.910.910.9), respectively.
@@
Figure 2 shows the results of the sensitivity analysis performed. The extended CIR estimates are substantially more sensitive to violations of conditional independence at the earliest time point, 30 days after the positive test, compared to 60 and 90 days. At 30 days, under negative dependence, meaning that those with longer symptom resolution times tend to respond more quickly to the survey, the estimated proportion with persistent symptoms is lower than in the primary analysis. Under positive dependence, the opposite pattern emerges and also reveals greater sensitivity. At 60 and 90 days, negative dependence has little discernible impact on the estimated proportion with ongoing symptoms, whereas positive dependence results in a moderately higher estimated proportion, particularly at 60 days.
Figure 2: Sensitivity analysis for deviation from conditionally uninformative response time assumption. From top to bottom, the rows correspond to the estimated proportion with ongoing symptoms at $30$, $60$, and $90$ days after the positive test. The $x$-axis represents the value of Kendall’s rank correlation coefficient $\tau$ used to encode conditional dependence in each sensitivity analysis. At $\tau=0$, in black, we give the estimate and confidence interval obtained from the proposed extended CIR procedure assuming conditional independence (primary analysis). Gray crosses represent estimates corresponding to different values of $\tau$.
Table 4 contains the results of the Cox proportional hazards regression, with hazard ratios larger than $1$ indicating faster symptom resolution. Participants identifying as male tended to experience faster symptom resolution, with an estimated hazard ratio for resolution of $1.21$ (1.041.041.041.041.421.421.421.42, $P=0.014$). Conversely, participants reporting fatigue at baseline tended to have longer lasting symptoms, with an estimated hazard ratio of $0.77$ (0.650.650.650.650.930.930.930.93, $P=0.005$). In terms of viral load, participants with Ct values under $30$ and participants who tested positive via rapid antigen tests had more persistent symptoms compared to participants with Ct values over $30$, with estimated hazard ratios of $0.78$ (0.640.640.640.640.960.960.960.96) and $0.76$ (0.600.600.600.600.960.960.960.96), respectively ($P=0.034$).
@@
        26–40 0.923 (0.767, 1.111)
        41–60 0.914 (0.749, 1.115)
        61 and over 1.075 (0.777, 1.487)
Male 1.214 (1.040, 1.417) 0.014
Symptoms, upper respiratory 1.085 (0.891, 1.322) 0.418
Symptoms, systemic 0.964 (0.802, 1.160) 0.699
Symptoms, gastrointestinal 0.885 (0.674, 1.164) 0.383
Symptoms, fatigue 0.774 (0.646, 0.927) 0.005
Comorbidities, cardiometabolic 0.809 (0.602, 1.087) 0.159
Comorbidities, respiratory 0.869 (0.690, 1.095) 0.234
Comorbidities, immunological 0.840 (0.490, 1.439) 0.526
Comorbidities, seasonal allergy 0.864 (0.741, 1.006) 0.060
Time since last vaccination 0.798
@@
         3–6 months 0.946 (0.746, 1.200)
         More than 6 months 0.931 (0.752, 1.151)
Viral load 0.034
@@
         PCR, under 30 Ct 0.782 (0.636, 0.962)
         Rapid antigen test 0.755 (0.596, 0.956)
@@ Discussion
A retrospective study in a comparable university population found that $36$% of survey respondents reported symptoms lasting at least 28 days from the end of the recommended $10$-day isolation period (Landry et al.,, 2023). Our results provide further evidence that, even in a predominantly young and vaccinated population, a substantial proportion of infected individuals are affected by persistent COVID-19 symptoms. Notably, because participants were asked only about persistent symptoms and not about new symptoms arising in the post-acute phase, our analysis may still represent an underestimate of the frequency of PASC, which may also be characterized by symptoms that differ from initial acute symptoms and arise after their resolution.
We also investigated potential risk factors for persistent symptoms using Cox proportional hazards regression, and found evidence that female sex, fatigue during the acute phase, and higher semi-quantitative viral load at time of testing are associated with slower resolution of symptoms. This is in line with previous studies that have identified female sex (Su et al.,, 2022), severe fatigue during the acute infection phase (Shen et al.,, 2023), and higher peak semi-quantitative viral load in the four weeks following diagnosis (Lu et al.,, 2023) as potential risk factors for PASC.
The validity of extended CIR hinges on the untestable assumption that survey response times and event indicators are conditionally independent given measured baseline covariates, which is more likely to be satisfied through the inclusion of potential risk factors as covariates in the analysis. If the survey response time depends on symptom duration even within strata defined by baseline covariates, the extended CIR procedure could be susceptible to bias. We derived one approach for formal sensitivity analysis in this setting, which allows investigators to examine changes in the estimated survival function as the dependence between response times and symptom resolution times, given measured covariates, deviates from zero. The development of additional sensitivity analysis methods based on alternative deviations is an important avenue of future research. Further work is also needed to better understand the practical implications of the poor boundary performance of CIR in order to provide a principled approach to choose appropriate values of $t_{0},t_{1}$ and $c_{0}$.
@@ Acknowledgements
We denote by $T$ the unobserved symptom duration after testing positive for SARS-CoV-2. For each study participant, we observe a vector $W\in\mathbb{R}^{p}$ of baseline covariates. In the context of current status sampling, the ideal data unit — that is, the data unit that would be observed given unlimited follow-up — consists of $W$, the survey response time $Y_{*}\in[b_{0},\infty)$, and the event indicator $\Delta_{*}\in\{0,1\}$. We denote by $c_{0}$ the maximum follow-up time of the study. In reality, we observe $Y:=\min(Y_{*},c_{0})$ and $\Delta:=\mathbbm{1}(Y_{*}<c_{0})\Delta_{*}$. For individuals who have not responded to the survey by time $c_{0}$, we do not observe the event indicator. Thus, the observed data unit is $Z:=(W,Y,\Delta)$. In this setting, survey non-response corresponds to having failed to respond to the survey prior to time $c_{0}$. The available data consist of $n$ independent observations $Z_{1},Z_{2},\ldots,Z_{n}$ draw from distribution $P_{0}$, which is assumed to lie in a nonparametric model $\mathcal{M}$. Here and after, we use the subscript 0 to denote a functional of the true data-generating distribution $P_{0}$, e.g., $E_{0}\{f(W)\}:=E_{P_{0}}\{f(W)\}$. For simplicity of notation, in what follows we consider estimation of the cumulative distribution function $t\mapsto\psi_{0}(t):=P_{0}(T\leq t)$. (In the main text, $\psi$ was used to denote the survival function, which is simply equal to one minus the distribution function).
@@
Without making strong modeling assumptions, the target regression function $\theta_{0}(y_{0})$ is not pathwise differentiable and is therefore not estimable at $n^{-\frac{1}{2}}$–rate. However, it is known to be monotone as a function of $y_{0}$ and can be viewed as the target of an isotonic regression procedure. Using isotonic regression to estimate $\theta_{0}(y_{0})$ allows for the application of established statistical theory (Westling and Carone,, 2020) to characterize the large-sample properties of the resulting estimator. We outline this isotonic regression procedure here.
For a generic $P\in\mathcal{M}$, we first define the outcome regression function $\mu_{P}:(y,w)\mapsto E_{P}(\Delta\,|\,W=w,Y=y)$ and density ratio $g_{P}:(y,w)\mapsto\pi_{P}(y\,|\,w)/f_{P}(y)$ with $\pi_{P}$ the conditional density of $Y$ given $W$ and $f_{P}$ the marginal density of $Y$. These nuisance functions play a key role in accounting for the possible dependence between $\Delta_{*}$ and $Y_{*}$. We use $F_{P}$ to denote the marginal distribution function of $Y$. We use $F_{P}$ as a domain transformation function in the CIR procedure. While the CIR procedure could be performed without the use of this domain transformation, its use results in a procedure that is invariant to strictly increasing transformations of $Y$ (Westling et al.,, 2020). In other words, changing the scale on which response times are measured will have no impact on the estimated survival function.
@@
with $Q_{P}$ the marginal distribution function of $W$ under sampling from $P$. Unlike $\theta_{0}(y_{0})$, the parameter $\Gamma_{0}(y_{0})$ is pathwise differentiable relative to the nonparametric model $\mathcal{M}$ and can therefore be estimated at $n^{-\frac{1}{2}}$–rate (Pfanzagl and Pfanzagl,, 1990). Per the results of Westling et al., (2020), the efficient influence function of $P\mapsto\Gamma_{P}(y_{0})$ relative to $\mathcal{M}$ evaluated at $P_{0}$ is given by
@@
Knowledge of this efficient influence functions enables standard debiasing approaches, such as the one-step correction (Pfanzagl,, 1982), to be used in estimation of $\Gamma_{0}(y_{0})$. This yields a debiased, asymptotically linear estimator $\Gamma_{n}(y_{0})$ of $\Gamma_{0}(y_{0})$, which is used in the isotonic regression procedure to construct an estimator of $\theta_{0}(y_{0})$. Our proposed procedure is as follows.
@@
    Let $\mathcal{R}_{n}$ denote the set of unique values of $Y_{1},Y_{2},\ldots,Y_{n}$. For each $y\in\mathcal{R}_{n}$ satisfying $t_{0}\leq y\leq t_{1}$, compute the one-step debiased estimator $\Gamma_{n}(y)$ of $\Gamma_{0}(y)$, given by $\Gamma_{n}(y):=\Gamma_{P_{n}}(y_{0})+\frac{1}{n}\sum_{i=1}^{n}D^{*}_{P_{n},y}(Z_{i})$.
@@
    Compute $\theta_{n}(y_{0}):=\psi_{n}\{F_{n}(y_{0})\}$ as an estimator of $\theta_{0}(y_{0})$.
We note that, were we not to apply the domain transformation $F_{P}$, we could perform a similar procedure using the primitive parameter $y_{0}\mapsto\iint\mathbbm{1}(u\leq y_{0})\mu_{P}(u,w)Q_{P}(dw)$. In the absence of survey nonresponse, this procedure would take the same form as the isotonization method of van der Vaart and van der Laan, (2006). However, it is not clear whether this approach can be framed in terms of isotonic regression using pseudo-outcomes, as described in the main text.
If for every point $y_{0}$ in the estimation domain $[t_{0},t_{1}]$, either $\mu_{n}$ converges to $\mu_{0}$ or $g_{n}$ converges to $g_{0}$, Theorem 1 of Westling et al., (2020) establishes that $\theta_{n}(y_{0})$ converges in probability to $\theta_{0}(y_{0})$ for each $y_{0}\in[t_{0},t_{1}]$. The conditions used to establish distributional convergence are more complex. Roughly speaking, we require that the product of the convergence rates of $\mu_{n}$ and $g_{n}$ to their respective targets is $n^{-\frac{1}{3}}$. If these conditions are satisfied, Theorem 2 of Westling et al., (2020) establishes that, for any $y_{0}\in(b_{0},c_{0})$ with $F_{0}(y_{0})\in(0,1)$, $n^{\frac{1}{3}}\left\{\theta_{n}(y_{0})-\theta_{0}(y_{0})\right\}$ converges in distribution to $\tau_{0}(y_{0})^{\frac{1}{3}}\mathbb{W}$, where $\mathbb{W}$ follows a standard Chernoff distribution (Groeneboom and Wellner,, 2001) and $\tau_{0}(y_{0}):=4\theta_{0}^{\prime}(y_{0})\kappa_{0}(y_{0})/f_{0}(y_{0})$ with